OpenStudy (anonymous):

put these equations in simplest form: y = (x-2)^2 + 4 4x + 2y = 14

OpenStudy (anonymous):

first one y=x^2-4x+4+4 so y=x^2-4x+8 second devide 2 each part 2x+y=7

OpenStudy (anonymous):

thank you :)